OpenStudy (anonymous):

Solve for X: 3e^(4x) = 6

OpenStudy (mathteacher1729):

How far have you gotten on this problem before getting stuck?

OpenStudy (anonymous):

Ln e^(4x^3) = Ln 6

OpenStudy (anonymous):

ln(2)/4

OpenStudy (anonymous):

4x^3 = Ln 6 dont know if its right

OpenStudy (anonymous):

thats right so far, keep going :)

OpenStudy (anonymous):

do i divide 4 and then cube

OpenStudy (mathteacher1729):

I would recommend starting by dividing both sides by 3. :)

OpenStudy (anonymous):

divide by four is correct :)

OpenStudy (anonymous):

that 3 is a power, it looks like this: \[4x^{3} = \ln(6) \Rightarrow x^{3} = \frac{\ln(6)}{4}\]

OpenStudy (anonymous):

so now we gotta get rid of that 3rd power above the x, so im going to take the cube root of both sides: \[x^{3} = \frac{\ln(6)}{4} \Rightarrow x = \sqrt[3]{\frac{\ln(6)}{4}}\] that would be your final answer, i dont have a calculator though, so i cant give you the exact number.

OpenStudy (anonymous):

that gives me 0.77 which is wrong though

OpenStudy (anonymous):

i think math teacher is right i get .173

OpenStudy (anonymous):

ti 89 type in press solve

OpenStudy (anonymous):

oh oh, i didnt start from the original problem, i started from the line 4x^3 = ln6. i didnt see the 3 in front of the e, sry >.<

OpenStudy (anonymous):

thats ok thanks!

OpenStudy (anonymous):

i have a ti 89 and i dont know what you mean MechE_MST